# references.bib

### v1

@@ 2.1 Preliminary definitions for p=2
[wang2022explicit] As an $\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}$-module, $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}}$ is freely generated by $1$
Note that by the classification of indecomposables in Proposition 2.3, every extension of $\mathrm{std}$ and every extension of a module by $\mathrm{std}$ splits. Thus $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is a direct summand of $Q_{m}(3,\mathbf{F}_{2})$ (whose complement is $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$), and we mainly consider $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$. $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is generated as a $\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}$-module by homogeneous copies of $\mathrm{std}$, so following [wang2022explicit], we consider generating representations of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ as homogeneous copies of $\mathrm{std}$ in a generators and relations presentation of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ with a minimal generator set.
2.1.1 Quasi-invariants at half-integers
Note that if $\mathbbm{k}$ is a field with char $\text{char\,}\mathbbm{k}\neq 2$ and $m\in\mathbb{Z}_{\geq 0}$, then for any $K\in\mathbbm{k}[x_{1},\dots,x_{n}]$, $(x_{i}-x_{j})^{2m}|(1-s_{ij})K$ implies $(x_{i}-x_{j})^{2m+1}|(1-s_{ij})K$ since $(1-s_{ij})K$ is $s_{ij}$-antiinvariant, hence the exponent $12m+$ 1 in the definition of quasi-invariant polynomials. But this does not hold in characteristic 2, since there is no concept of antiinvariants. Indeed, one can check that for $K=x_{1}^{2}+x_{2}^{2}$, we have $(x_{i}-x_{j})^{2}|(1-s_{ij})K$ for all $i,j$, but $(x_{i}-x_{j})^{3}\nmid|(1-s_{ij})K$ if $i=1,2$, $j\neq 1,2$.
2.2 Preliminary definitions for $p=3$
Proposition 2.7 ([Alperin_1986]).
  1. 1)
  2. 2)
  3. 3)
  4. 4)
  5. 5)
  6. 6)
Example 2.8.
The space $V\subset\mathbf{F}_{3}[x_{1},x_{2},x_{3}]$ consisting of homogeneous linear polynomials is a copy of $\mathrm{triv-sign-triv}$. Indeed, the space $T\subset V$ spanned by $x_{1}+x_{2}+x_{3}$ is a copy of $\mathrm{triv}$. One can check $x_{1}-x_{2}\in V/T$ is acted by negation by all transpositions in $S_{3},$ so $x_{1}-x_{2}$ generates a copy of $\mathrm{sign}$ in $V/T$. Let $W$ be spanned by $x_{1}-x_{2}$ and $x_{1}+x_{2}+x_{3}$. Then $V/W$ is one dimensional, and one can check that it is a copy of $\mathrm{triv}$. Finally, it is easy to show that there are no copies of $\mathrm{triv}$ or $\mathrm{sign}$ in $V$ other than $T$. In particular, since $V$ has a unique irreducible subrepresentation, it is indecomposable, and we conclude that it is a copy of $\mathrm{triv-sign-triv}$.
Similarly to the $p=2$ case, we define $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ and $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{triv}}$ to be the direct sum of all copies of $\mathrm{sign}$ and $\mathrm{triv}$ in $Q_{m}(3,\mathbf{F}_{3})$ respectively.
Proposition 2.9 ([wang2022explicit]).
As $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$-modules,
  1. 1)
    $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{triv}}$ is freely generated by $1$.
  2. 2)
    $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ is freely generated by $\prod_{i<j}(x_{i}-x_{j})^{2m+1}$.
Next we define $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ as the direct sum of all copies of $\mathrm{sign},$ $\mathrm{triv},$ and $\mathrm{sign-triv}.$ For this paper we consider generators of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ to be homogeneous polynomials other than $1$ and $\prod_{i<j}(x_{i}-x_{j})^{2m+1}$ such that they are in the $(-1)$-eigenspace of $s_{12}$ and are in a generators and relations presentation of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ as an $F_{3}[x_{1},x_{2},x_{3}]-$module with the least number of generators. Moreover, if $K$ is a generator of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ then it necessarily generates a copy of $\mathrm{sign-triv}$ since we assumed $K$ neither generates $\mathrm{triv}$ nor $\mathrm{sign}.$
Remark 2.10.
Similar to in the $p=2$ case, we cannot define $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ to exclude copies of $\mathrm{sign}$ since we can add elements of $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign}}$ to copies of $\mathrm{sign-triv}$ and still obtain a copy of $\mathrm{sign-triv}.$ For example, the spaces spanned by $(x_{1}^{6}-x_{2}^{6})(x_{1}+x_{2}+x_{3})^{3}$, $(x_{1}^{6}+x_{2}^{6}+x_{3}^{6})(x_{1}+x_{2}+x_{3})^{3}$ and $\prod_{i<j}(x_{i}-x_{j})^{3}+(x_{1}^{6}-x_{2}^{6})(x_{1}+x_{2}+x_{3})^{3}$, $(x_{1}^{6}+x_{2}^{6}+x_{3}^{6})(x_{1}+x_{2}+x_{3})^{3}$ generate two copies of $\mathrm{sign-triv}$ in $Q_{1}(3,\mathbf{F}_{3})$, and their sum contains $\prod_{i<j}(x_{i}-x_{j})^{3}\in Q_{1}(3,\mathbf{F}_{3})_{\mathrm{sign}}$.
Remark 2.11.
One could define subspaces of $Q_{m}(3,\mathbf{F}_{3})$ for $\mathrm{triv-sign-triv},\mathrm{sign-triv-sign},$ $\mathrm{triv-sign}$ similar to $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$, however this is not particularly helpful, as unlike in $p=2$, we cannot decompose $Q_{m}(3,\mathbf{F}_{3})$ into a direct sum of subspaces of this form. $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ is still relevant, as it is the critical piece to understanding quasi-invariants in characteristic 3, as we see in Sections 4 and 5.
3 Quasi-invariants in characteristic 2
In this section we write down explicit generators for $Q_{m}(3,\mathbf{F}_{2})$ and prove Theorem 1.3 for $p=2$. Note that we already know the structure of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}}$ from Proposition 2.6. We start by extending this to $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$.
Proposition 3.1.
As a $\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}$-module, $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}-\mathrm{triv}}$ is freely generated by $1$ and $E_{\mathrm{triv-triv}}\prod(x_{i}-x_{j})^{2m}$.
@@ 3 Quasi-invariants in characteristic 2
Let $K$ be a nonsymmetric element of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}-\mathrm{triv}}$ so that $(x_{i}+x_{j})^{2m+1}|(1+s_{ij})K.$ Because
@@
$(1+s_{ij})K=P\prod(x_{i}-x_{j})^{2m+1}$ for some symmetric polynomial $P.$ Letting $G=E_{\mathrm{triv-triv}}\prod(x_{i}-x_{j})^{2m}$ yields $(1+s_{ij})G=\prod(x_{i}-x_{j})^{2m+1}.$ Thus $(1+s_{ij})PG=(1+s_{ij})K$ and $(1+s_{ij})(PG-K)=0,$ so $PG-K$ is symmetric and $K$ is generated by $G$ and $1$ Moreover since $G$ is not symmetric, $P$ and $G$ have no relation implying freeness. ∎
We have an explicit description of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$, so it remains to compute the generators and relations of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$. A number of the properties of $Q_{m}(3,\mathbf{F}_{p})$ for $p>3$ found in [wang2022explicit] are true for $Q_{m}(3,\mathbf{F}_{2}).$ We prove these first.
If $V$ is a copy of $\mathrm{std}$, then we denote by $V_{ij}$ the $1$-eigenspace of $s_{ij}$ in $V$.
@@
For the first statement, $K+sK+s^{2}K=0$ holds for any copy of $\mathrm{std}.$ For the next, suppose $\{i,j,l\}=\{1,2,3\}$ for some integer $l.$ Then $(1-s_{il})K=s_{jl}K,$ so $(x_{i}-x_{l})^{2m+1}|s_{jl}K$, implying $(x_{i}-x_{j})^{2m+1}|K.$ The second statement follows from the proof in [wang2022explicit]. ∎
@@
Let $V$ be a generating representation of $Q_{m}(3,\mathbbm{k})_{\mathrm{std}}$ and let $K\in V_{ij}$. Let us write $K=(x_{i}-x_{j})^{2m+1}K^{\prime}$. Then $K^{\prime}$ is not divisible by any nonconstant symmetric polynomial.
The proof of this statement is identical to the one in [wang2022explicit].
@@
Let $V,W$ be distinct generating representations of $Q_{m}(3,\mathbbm{k})_{\mathrm{std}}$. Let $K\in V_{12},L\in W_{12}$. Then $KL+s_{13}Ks_{23}L$ is a nonsymmetric element of $Q_{m}(3,\mathbbm{k})_{\mathrm{triv-triv}}$ and we have $\deg V+\deg W\geq 6m+3$.
@@
$KL+s_{13}Ks_{23}L$ is an element of $Q_{m}(3,\mathbf{F}_{2})$ since the quasi-invariants form a ring by Proposition 2.1. We have that
@@
Using that $K$ and $L$ lie in copies of $\mathrm{std},$ one can check that these are each the same symmetric polynomial. Thus $KL+s_{13}Ks_{23}L$ lies in a quotient of a copy of $\mathrm{triv-triv}$. Note that by the same argument as in [wang2022explicit], we have $Ks_{23}L+s_{23}KL\neq 0$, so $KL+s_{13}Ks_{23}L$ is nonsymmetric and must generate a copy of $\mathrm{triv-triv}$.
By Proposition 2.6, $KL+s_{13}Ks_{23}L$ has degree at least $6m+3,$ so $\deg V+\deg W\geq 6m+3$ as desired. ∎
@@
$(1+s_{12})(KT+s_{13}Ks_{23}T)=s_{23}Ks_{13}T+s_{13}Ks_{23}T=Q\prod(x_{i}-x_{j})^{2m+1}$
for some symmetric polynomial $Q.$ From there we may proceed identically to [wang2022explicit]. ∎
@@
Let $a$ be the largest natural number such that $12^{a}<2m+$ 1 . Then $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is freely generated by $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2m+1-2^{a}}.$
@@
Note that when $m$ is an integer, the degrees of the generators in this theorem agree with the degrees conjectured in [Ren_2020]. In particular, when $2^{a+1}$ is one of $23m+1,3m+$ 2, we actually have that the Hilbert series of $Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{Q})$ agree, so $(x_{1}-x_{2})^{2^{a+1}}$, $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2m+1-2^{a}}$ are the reductions modulo 2 of the generators of $Q_{m}(3,\mathbf{Q})$, when written as integer polynomials with coprime coefficients.
@@
Let $j$ be a half-integer, and suppose that $Q_{j-\frac{1}{2}}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is freely generated by $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2j-2^{a}}$ where $2^{a}$ is the greatest such power of $2$ less than $2j.$ If $2j\neq 2^{a+1},$ then $2^{a}$ is the largest power of $2$ less than $2j+1,$ so $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2j+1-2^{a}}$ are both in $Q_{j}(3,\mathbf{F}_{2}).$ Further, $(x_{1}-x_{2})^{2^{a+1}}$ must be a generator and if $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2j+1-2^{a}}$ is a not a generator, by Lemma 3.4, $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2j+1-2^{a}}$ is generated by $(x_{1}-x_{2})^{2^{a+1}}$ which implies a relation between $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2j-2^{a}}.$ Because they freely generate $Q_{j-\frac{1}{2}}(3,\mathbf{F}_{2}),$ this is impossible. Thus $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i}-x_{j})^{2j+1-2^{a}}$ freely generate $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ by Lemma 3.5.
If $2j+1=2^{a+1},$ then both $(x_{2}-x_{3})^{2^{a+1}}$ and $(x_{2}-x_{3})^{2^{a+2}}$ lie in $Q_{j}(3,\mathbf{F}_{2}).$ The former is a generator by our inductive hypothesis. Since $2^{a+1}+2^{a+2}=6j+3,$ if the latter is not a generator, then by Lemma 3.4, $(x_{2}-x_{3})^{2^{a+2}}$ is generated by $(x_{2}-x_{3})^{2^{a+1}}$, which is false. Thus $(x_{2}-x_{3})^{2^{a+1}}$ and $(x_{2}-x_{3})^{2^{a+2}}$ freely generate $Q_{j}(3,\mathbf{F}_{2})$ by Lemma 3.5 as desired.
4 Properties of $3$ variable quasi-invariants
Similar to in the $p=2$ case, we can adapt many of the properties of $Q_{m}(3,\mathbf{F}_{p})$ for $p>3$ found in [wang2022explicit] to the $p=3$ case. We accomplish this by converting $\mathrm{std}$ to $\mathrm{sign-triv}$. For example, in $Q_{0}(3,\mathbf{F}_{3})$, the space spanned by $x_{1}-x_{2},x_{1}-x_{3}$ is a copy of $\mathrm{std}$. However, in $Q_{0}(3,\mathbf{F}_{3})$, the space spanned by $x_{1}-x_{2},x_{1}-x_{3}$ becomes a copy of $\mathrm{sign-triv}$. Using this we may show that there are equivalents of Lemma 3.2 to Lemma 3.5 from [wang2022explicit] in characteristic $3$.
We define $V_{ij}^{-}$ to be the $(-1)$-eigenspace of $s_{ij}$ in $V$, where $V$ is a copy of $\mathrm{std}$ or $\mathrm{sign-triv}$. Note that if $v\in V_{ij}^{-}$ we have $v=s_{23}v+s_{13}v$. The following lemma and corollary correspond to Lemma 3.2 and Corollary 3.3 respectively from [wang2022explicit].

### v2

@@ 3 Quasi-invariants in characteristic 2
As an $\mathbf{F}_{2}[x_{1},x_{2},x_{3}]^{S_{3}}$-module, $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}-\mathrm{triv}}$ is freely generated by $1$ and $E_{\mathrm{triv-triv}}\prod(x_{i_{1}}-x_{i_{2}})^{2m}$.
@@
Let $K$ be a nonsymmetric element of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv}-\mathrm{triv}}$ so that $(x_{i_{1}}-x_{i_{2}})^{2m+1}$ divides $(1+s_{i_{1}i_{2}})K$. Because
@@
we have $(1+s_{i_{1}i_{2}})K=P\prod(x_{i_{1}}-x_{i_{2}})^{2m+1}$ for some symmetric polynomial $P$. Letting $G=E_{\mathrm{triv-triv}}\prod(x_{i_{1}}-x_{i_{2}})^{2m}$ yields $(1+s_{i_{1}i_{2}})G=\prod(x_{i_{1}}-x_{i_{2}})^{2m+1}$. Thus $(1+s_{i_{1}i_{2}})PG=(1+s_{i_{1}i_{2}})K$ and $(1+s_{i_{1}i_{2}})(PG-K)=0$, so $PG-K$ is symmetric and $K$ is generated by $G$ and $1$. Moreover, since $G$ is not symmetric, $P$ and $G$ have no relation implying freeness. ∎
We have an explicit description of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{triv-triv}}$, so it remains to compute the generators and relations of $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$. A number of the properties of $Q_{m}(3,\mathbf{F}_{p})$ for $p>3$ found in [13] are true for $Q_{m}(3,\mathbf{F}_{2})$. We prove these first.
If $V$ is a copy of $\mathrm{std}$, then we denote by $V_{i_{1}i_{2}}$ the $1$-eigenspace of $s_{i_{1}i_{2}}$ in $V$.
@@
For the first statement, $K+sK+s^{2}K=0$ holds for any copy of $\mathrm{std}$. For the next, suppose $\{i_{1},i_{2},i_{3}\}=\{1,2,3\}$ for some integer $i_{3}$. Then $(1-s_{i_{1}i_{3}})K=s_{i_{2}i_{3}}K$, so $(x_{i_{1}}-x_{i_{3}})^{2m+1}|s_{i_{2}i_{3}}K$, implying $(x_{i_{1}}-x_{i_{2}})^{2m+1}|K$. The second statement follows from the proof in [13]. ∎
@@
Let $V$ be a generating representation of $Q_{m}(3,\mathbbm{k})_{\mathrm{std}}$ and let $K\in V_{i_{1}i_{2}}$. Let us write $K=(x_{i_{1}}-x_{i_{2}})^{2m+1}K^{\prime}$. Then $K^{\prime}$ is not divisible by any nonconstant symmetric polynomial.
The proof of this statement is identical to the one in [13].
@@
Let $V$, $W$ be distinct generating representations of $Q_{m}(3,\mathbbm{k})_{\mathrm{std}}$. Let ${K\in V_{12}}$, ${L\in W_{12}}$. For $\sigma K\sigma L:=(\sigma K)(\sigma L)$, we have that $KL+s_{13}Ks_{23}L$ is a nonsymmetric element of $Q_{m}(3,\mathbbm{k})_{\mathrm{triv-triv}}$ and $\deg V+\deg W\geq 6m+3$.
@@
$KL+s_{13}Ks_{23}L$ is an element of $Q_{m}(3,\mathbf{F}_{2})$ since the quasi-invariants form a ring by Proposition 2.1. Using that $s_{12}K=K$ and $s_{12}L=L$, we have that
@@
One can check that each polynomial is a transposition of another and that they are symmetric due to the structure of $\mathrm{triv-triv}$, so they are all the same symmetric polynomial. Thus ${KL+s_{13}Ks_{23}L}$ lies in a quotient of a copy of $\mathrm{triv-triv}$. Note that by the same argument as in [13], we have $Ks_{23}L+s_{23}KL\neq 0$, so $KL+s_{13}Ks_{23}L$ is nonsymmetric and must generate a copy of $\mathrm{triv-triv}$.
By Proposition 3.1, $KL+s_{13}Ks_{23}L$ has degree at least $36m+$ 3, so $\deg V+\deg W\geq 6m+3$ as desired. ∎
@@
$(1+s_{12})(KT+s_{13}Ks_{23}T)=s_{23}Ks_{13}T+s_{13}Ks_{23}T=Q\prod(x_{i_{1}}-x_{i_{2}})^{2m+1}$
for some symmetric polynomial $Q$. From there we may proceed identically to [13]. ∎
@@
Let $a$ be the largest natural number such that $12^{a}<2m+$ 1. Then $Q_{m}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is freely generated by $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2m+1-2^{a}}$.
@@
Note that when $m$ is an integer, the degrees of the generators in this theorem agree with the degrees conjectured in [11]. In particular, when $2^{a+1}$ is one of $13m+$ 1, ${3m+2}$, we actually have that the Hilbert series of $Q_{m}(3,\mathbf{F}_{2})$ and $Q_{m}(3,\mathbf{Q})$ agree, so $(x_{1}-x_{2})^{2^{a+1}}$, $(x_{1}-x_{2})^{2^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2m+1-2^{a}}$ are the reductions modulo 2 of the generators of $Q_{m}(3,\mathbf{Q})$, when written as integer polynomials with coprime coefficients.
@@
Let $j$ be a half-integer, and suppose that $Q_{j-\frac{1}{2}}(3,\mathbf{F}_{2})_{\mathrm{std}}$ is freely generated by $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2j-2^{a}}$, where $2^{a}$ is the greatest such power of $2$ less than $2j$. If ${2j\neq 2^{a+1}}$, then $2^{a}$ is the largest power of $2$ less than $12j+$ 1, so $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2j+1-2^{a}}$ are both in $Q_{j}(3,\mathbf{F}_{2})$. Further, $(x_{1}-x_{2})^{2^{a+1}}$ must be a generator and if $(x_{1}-x_{2})^{2^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2j+1-2^{a}}$ is a not a generator, by Lemma 3.4, $(x_{1}-x_{2})^{2^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2j+1-2^{a}}$ is generated by $(x_{1}-x_{2})^{2^{a+1}}$ which implies a relation between $(x_{1}-x_{2})^{2^{a+1}}$ and ${(x_{1}-x_{2})^{2^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2j-2^{a}}}$. Because they freely generate $Q_{j-\frac{1}{2}}(3,\mathbf{F}_{2})$, this is impossible. Thus $(x_{1}-x_{2})^{2^{a+1}}$ and $(x_{1}-x_{2})^{2^{a}}\smash{\prod(x_{i_{1}}-x_{i_{2}})^{2j+1-2^{a}}}$ freely generate $Q_{j}(3,\mathbf{F}_{2})_{\mathrm{std}}$ by Lemma 3.5.
If $2j=2^{a+1}$, then both $(x_{2}-x_{3})^{2^{a+1}}$ and $(x_{2}-x_{3})^{2^{a+2}}\prod(x_{i_{1}}-x_{i_{2}})^{2j+1-2^{a}}$ lie in $Q_{j}(3,\mathbf{F}_{2})_{\mathrm{std}}$. The former is a generator by our inductive hypothesis. Since $32^{a+1}+2^{a+2}+3=6j+$ 3, if the latter is not a generator, then by Lemma 3.4, $(x_{2}-x_{3})^{2^{a+2}\prod(x_{i_{1}}-x_{i_{2}})^{2j+1-2^{a}}}$ is generated by $(x_{2}-x_{3})^{2^{a+1}}$, which is false. Thus $(x_{2}-x_{3})^{2^{a+1}}$ and $(x_{2}-x_{3})^{2^{a+2}}\prod(x_{i_{1}}-x_{i_{2}})^{2j+1-2^{a}}$ freely generate $Q_{j}(3,\mathbf{F}_{2})_{\mathrm{std}}$ by Lemma 3.5 as desired. ∎
4 Properties of $\boldsymbol{3}$ variable quasi-invariants
Similarly to the $p=2$ case, we can adapt many of the properties of $Q_{m}(3,\mathbf{F}_{p})$ for $p>3$ found in [13] to the $p=3$ case. We accomplish this by converting $\mathrm{std}$ to $\mathrm{sign-triv}$. For example, in $Q_{0}(3,\mathbf{F}_{p})$ for $p>3$, the space spanned by $x_{1}-x_{2}$, $x_{1}-x_{3}$ is a copy of $\mathrm{std}$. However, in $Q_{0}(3,\mathbf{F}_{3})$, the space spanned by $x_{1}-x_{2}$, $x_{1}-x_{3}$ becomes a copy of $\mathrm{sign-triv}$. Using this, we may show that there are equivalents of Lemmas 3.2–3.5 from [13] in characteristic $3$.
We define $V_{i_{1}i_{2}}^{-}$ to be the $(-1)$-eigenspace of $s_{i_{1}i_{2}}$ in $V$, where $V$ is a copy of $\mathrm{std}$ or $\mathrm{sign-triv}$. Note that if $v\in V_{i_{1}i_{2}}^{-}$ we have $v=s_{23}v+s_{13}v$. The following lemma and corollary correspond to Lemma 3.2 and Corollary 3.3 from [13], respectively.